# RepairLLaMA: Efficient Representations and Fine-Tuned Adapters for Program Repair

### v3

@@ 3.1. Research Questions
  • RQ3 (RepairLLaMA vs. ChatGPT-based APR) How does RepairLLaMA compare against state-of-the-art ChatGPT-based program repair?
3.2. Implementation
@@
Model to Fine-Tune Per the criteria of subsection 2.3, we choose CodeLlama-7b (Roziere et al., 2023) as our initial LLM. CodeLLaMA is a publicly available LLM released in 2023 and is trained on 500B code tokens. Per the experiments reported in (Roziere et al., 2023), CodeLLaMA outperforms GPT-3.5 on two code generation benchmarks.
@@
Fine-tuning Dataset We choose Megadiff (Monperrus et al., 2021) as the fine-tuning dataset, and process all samples into the different code representations. First, the function pairs – each comprising a buggy version and its fixed counterpart – are extracted along with their corresponding diff identifiers. Subsequently, we eliminate pairs that do not change single functions, and remove duplicate pairs through textual comparison. After that, we compute our custom code representations. We keep only samples whose total length (input plus output) is shorter than 1024 tokens measured by the LLM tokenizer. Consequently, the fine-tuning datasets range from 30,000 to 50,000 fine-tuning pairs (see our appendix repository).
@@
Evaluation Benchmark We select two Java benchmarks for our evaluation: Defects4J (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023). Following recent related work (Jiang et al., 2021b; Xia et al., 2022; Xia and Zhang, 2023), we scope our evaluation to single-function bugs. Defects4J comprises 835 real-world bugs from 17 open-source Java projects, from which we identify 488 single-function bugs. HumanEval-Java is a bug benchmark containing artificial bugs inserted in HumanEval (Chen et al., 2021b) Java programs. Contrary to Defects4J, HumanEval-Java suffers from less data leakage in the pre-training data since it is much more recent than Defects4J. HumanEval-Java contains 164 single-function bugs.
@@
Fine-Tuning We fine-tune CodeLLaMA with LoRA for each of our curated code representations with the same hyper-parameter settings: we set the learning rate to 5e-4 with cosine decay, max input length to 1024, training epoch to 2, and batch size to 16 per GPU, and we use Adam_W as the optimizer. For LoRA, we use a rank of 8, alpha of 16, dropout of 0.05, and inject the adaptation matrices in the q_proj and v_proj layers. Using the same hyper-parameter settings for each code representation ensures fair comparison. Each fine-tuning run is executed on a server with 4xA100 40GB GPUs.
@@
Inference In inference, we employ beam search as our decoding strategy with a beam size of 10 per previous research (Jiang et al., 2023). Hence, for each bug, we generate 10 candidate patches. We use the HuggingFace transformers library to implement all fine-tuning and inference experiments. Inference is run on a single A100 40GB GPU.
3.3. Patch Assessment
@@
Following related work (Xia et al., 2022; Jiang et al., 2023; Ye and Monperrus, 2023), we compute the following repair effectiveness metrics. A plausible patch is defined as one that successfully passes all test cases. An exact match patch is textually identical to a developer-provided reference patch. To refine our evaluation, we further assess the syntactical equivalence between the generated patches and the reference patches, through AST match. This enables us to compute performance regardless of formatting and indentation changes. This process involves converting plausible and reference patches into abstract syntax trees (Pawlak et al., 2015) and subsequently utilizing AST differencing (Falleri et al., 2014) to compare their ASTs for discrepancies. A plausible patch with no AST differences compared to the reference patch is classified as an AST match patch. It is also more scalable and accurate than manually checking plausible patches for correctness without expertise in the programs under repair. For all three metrics, the higher the metric, the better the performance. We validate the candidate patches on a workstation with an 18-core Intel Core i9-10980XE CPU and 128 GB of RAM, operating under Ubuntu 22.04.3 LTS.
3.4. Methodology for RQ1
@@
The objective of RQ1 is to investigate the most effective code representations for fine-tuning an LLM for program repair. While existing research has delved into the utility of LLMs for program repair, the impact of the code representations, such as their authenticity, has been overlooked. Additionally, it is crucial to note that variations in code representations may yield substantial differences in performance for fine-tuned LLMs (Huang et al., 2023a). Consequently, in RQ1, we curate 6 realistic code representations motivated in 2.4 and execute a series of experiments to measure their performance. We fine-tune an LLM as described in 3.2. We prompt the model to generate 10 patches for each bug using beam search decoding. We then evaluate the generated patches as outlined in subsection 3.3, to measure the effectiveness of each code representation.
Baseline. Since CodeLLaMA performs powerful zero-shot learning ability in other code-related tasks (Roziere et al., 2023), we use the same method to prompt CodeLLaMA-7B with the infilling scheme (IR3xOR2) on our selected benchmark and take the result as the baseline, which can help us better evaluate the performance of every code representation and measure the effectiveness of fine-tuning.
3.5. Methodology for RQ2
@@
The objective of RQ2 is to evaluate the respective effectiveness of parameter-efficient and full-parameter fine-tuning. Generally, parameter-efficient fine-tuning methods represent a trade-off between computational cost and model performance, which allows us to train LLMs with limited computational resources. While traditional full-parameter fine-tuning approaches often yield better results, they come at the expense of significantly higher memory requirements and a large-scale fine-tuning dataset. In other words, fully fine-tuning an LLM on a small fine-tuning dataset may drop into the overfitting problem. We explore and compare the effectiveness of parameter-efficient and full-parameter fine-tuning in the specific context of program repair, which has never been done to the best of our knowledge.
@@
Baseline. We consider four baselines in RQ2. One is the baseline considered in RQ1. The second one comprises the full-parameter fine-tuned version of RepairLLaMA, a powerful but naive approach to fine-tuning a large language model for program repair. Here, we use the same hyper-parameters as in LoRA fine-tuning, apart from a lower learning rate of 2e-5. The third is Jiang et al.’s work (Jiang et al., 2023), where several LLMs are fully fine-tuned for program repair. We compare against the best performing model reported, the fine-tuned version of Incoder-6B (Fried et al., 2022). The last one is RAP-Gen (Wang et al., 2023b), the current state-of-the-art fine-tuned LLM for program repair. Its key novelty is to use relevant fix patterns retrieved from a codebase of previous bug-fix pairs to augment the buggy input, which guides the model to generate accurate patches by learning from historical repair examples.
3.6. Methodology for RQ3
Figure 4. The prompt used to prompt GPT-3.5 and GPT-4 as a strong baseline to generate patches.
@@ 3.6. Methodology for RQ3
The objective of RQ3 is to study how RepairLLaMA compares against state-of-the-art ChatGPT-based program repair. Recently, related work (Xia and Zhang, 2023; Zhang et al., 2023) has shown that GPT-3.5-Turbo and GPT-4 achieve state-of-the-art results on program repair. To this end, we compare RepairLLaMA in two experiments.
First, we zero-shot prompt gpt-3.5-turbo-0613 and gpt-4-0613 to generate 10 patches for each bug. The prompt is shown in Figure 4, which is built by integrating an effective prompt from Zhang et al.’s work and our curated best code representation. It instructs LLMs to generate the fixed code chunk to replace the ¡FILL_ME¿ token.
Second, we compare RepairLLaMA with ChatRepair on Defects4J v1.2, since ChatRepair’s authors mainly evaluate it on single-function bugs in Defects4 v1.2. To the best of our knowledge, ChatRepair achieves state-of-the-art performance on Defects4J, and such a comparison enables us to fully evaluate the effectiveness of RepairLLaMA. We utilize OpenAI’s official APIs to call gpt-3.5-turbo-0613 to conduct related experiments on Dec. 1, 2023.
4. Experimental Results
4.1. Results of RQ1 (Code Representations for Fine-Tuning)
@@
Table 2. Repair results of different code representations for fine-tuning an LLM for program repair. Our best model, RepairLLaMA using IR4xOR2, significantly improves over the baseline in both test benchmarks.
Code Representations Defects4J v2
@@
(164 bugs)
Plausible AST Match
Exact Match
Plausible AST Match Exact Match
IR3 x OR2 (baseline, no fine-tuning) 133 $71$
52
107 $81$ $71$
IR1 x OR1 79 $31$
29
78 $54$ $52$
IR1 x OR3 41 $17$
15
39 $21$ $21$
IR1 x OR4 12 $2$
2
5 $2$ $2$
IR2 x OR2 198 $122$
121
118 $77$ $69$
IR3 x OR2 154 $87$
84
103 $68$ $63$
IR4 x OR2 (RepairLLaMA) 195 125 124 118 82 75
@@ 5 $2$ $2$
In RQ1, we investigate the most effective code representations for fine-tuning an LLM for program repair. The results of the evaluation are presented in subsection 4.1, which shows the effectiveness of each code representation setting on both test benchmarks. The table is structured as follows: the first column displays the code representations used to fine-tune CodeLLaMA-7B, the second and third meta-columns show the repair effectiveness results on, respectively, Defects4J v2 (single-function) and HumanEval-Java (single-function). Recall that the repair effectiveness evaluation is measured by three metrics described in subsection 3.3.
@@
Our baseline uses the original non-fine-tuned CodeLLaMA-7B with the IR3xOR2 code representation. Recall that this is the most effective way to prompt the non-fine-tuned model as discussed in subsection 3.4. Our results show that it plausibly repairs 133 Defects4J v2 and 107 HumanEval-Java bugs. Moreover, it correctly repairs 52 Defects4J v2 and 71 HumanEval-Java bugs with patches that textually exactly match the developer-written ones. Furthermore, when considering AST match, it can repair 71 Defects4J v2 and 81 HumanEval-Java bugs with patches that are syntactically equal to the developer-written ones.
@@
The fine-tuned model’s effectiveness depends on the code representations. First, we observe that the three code representation pairs that do not have access to fault localization (IR1xORX) perform considerably worse than both the baseline and other code representations. These results show that fault localization signals are crucial for program repair. To that extent, all representations that simply use the full function as the input and ask to “fix the bug”, can be considered too naive. Our work demonstrates that tailoring code representations with fault localization is a necessary step in the context of program repair. No pre-training objective has access to fault-localization signals, which further validates the need for fine-tuning.
@@
Second, we observe that, within code representations that use fault localization signals, fine-tuned models significantly outperform the baseline on Defects4J v2 compared to the baseline (first row). For the same code representation as the baseline (first row), the fine-tuned model (sixth row, IR3xOR2) can plausibly repair 154 (+21) bugs, exactly repair 84 bugs (+32), and syntactically correctly repair 87 bugs (+16), respectively.
When considering the remaining two code representations (IR2xOR2, IR4xOR2), the fine-tuned models perform even better. The best model, RepairLLaMA, fine-tuned with the code representation IR4xOR2, plausibly repairs 195 (+62) Defects4J bugs than the baseline, exactly repairs 124 (+72), and syntactically correctly repair 125 bugs (+54), respectively. These results show that fine-tuning the initial LLM with repair-specific representations significantly improves the repair effectiveness over a non-fine-tuned model. Tailored code representations with bug localization signals allow the fine-tuned model to repair more bugs because these signals enable it to pay more attention to the buggy code.
@@
To strengthen the external validity of our analysis beyond Defects4J, we perform the same experiment on HumanEval-Java. On HumanEval-Java, RepairLLaMA also achieves better repair effectiveness than the baseline (75 vs 71 exact match and 82 vs 81 ast match), confirming the results observed on Defects4J.
Third, we discuss the alignment between the repair-specific code representation and the pre-training objectives. IR2 and IR4 are input representations that include the same signals: the original function, fault localization, and the original buggy code. However, we see a performance gap between both representations. IR4 uses the same infilling scheme during pre-training while providing bug location signals by including the original buggy code in comments. IR2, on the other hand, mimics a similar scheme but provides bug location signals through fault localization comments, which is further away from the pre-training distribution. This difference explains the differences between IR4xOR2 vs. IR2xOR2: the former has a lower representation gap with pre-training, thus leading to better repair effectiveness. This finding holds for both benchmarks (Defects4J and HumanEval-Java) and for both metrics (Exact Match and AST Match).
@@
Figure 5. Exact match patch generated by RepairLLaMA for Math-86 from Defects4J v2. In this multi-location bug, RepairLLaMA is able to fix two distant buggy locations.
@@
In addition to comparing performance, we explore the effectiveness of our single-chunk prompting approach in addressing multi-chunk bugs. Our findings reveal that RepairLLaMA correctly repairs 30 instances of such bugs. For example, RepairLLaMA correctly fixes a complex multi-chunk bug, Math-86, from the Defects4J v2, as illustrated in Figure 5. Math-86 presents two error sections that require simultaneous attention and correction: 1) the removal of an if block that throws an exception, and 2) the introduction of a new if condition. Note that these two sections have more than 20 lines of distance between each other, showing that RepairLLaMA can fix bugs where the multiple edit locations are far away from each other. To the best of our knowledge, RepairLLaMA is the first program repair approach to correctly fix Math-86.
@@
Figure 6. Exact match patch generated by RepairLLaMA for STRONGEST_EXTENSION from HumanEval-Java. In this multi-location bug, RepairLLaMA modifies two if blocks and encapsulates two other statements in a new if block.

### v5

@@ 3.1 Research Questions
  • RQ3 (RepairLLaMA vs. ChatGPT-based APR): How does RepairLLaMA compare against state-of-the-art ChatGPT-based program repair?
3.2 Implementation
@@
Model to Fine-Tune Per the criteria of subsection 2.3, we choose CodeLlama-7b [24] as our initial LLM. CodeLLaMA is a publicly available LLM released in 2023 and is trained on 500B code tokens. Per the experiments reported in [24], CodeLLaMA outperforms GPT-3.5 on two code generation benchmarks.
@@
Fine-tuning Dataset We choose Megadiff [33] as the fine-tuning dataset, and process all samples into the different code representations. First, the function pairs – each comprising a buggy version and its fixed counterpart – are extracted along with their corresponding diff identifiers. Subsequently, we eliminate pairs that do not change single functions, and remove duplicate pairs through textual comparison. After that, we compute our custom code representations. We keep only samples whose total length (input plus output) is shorter than 1024 tokens measured by the LLM tokenizer. Consequently, the fine-tuning datasets range from 30,000 to 50,000 fine-tuning pairs (see our appendix repository).
@@
Evaluation Benchmark We select three Java benchmarks for our evaluation: Defects4J [22], HumanEval-Java [8], and GitBug-Java [23]. Following recent related work [34, 11, 14], we scope our evaluation to single-function bugs, as defined in subsection 2.2. Defects4J comprises 835 real-world bugs from 17 open-source Java projects, from which we identify 488 single-function bugs. HumanEval-Java is a bug benchmark containing artificial bugs inserted in HumanEval [35] Java programs. HumanEval-Java contains 162 single-function bugs. GitBug-Java is a bug benchmark of recent bugs, collected from the 2023 commit history of 55 open-source repositories, comprising 90 single-function bugs. Contrary to Defects4J, GitBug-Java suffers from less data leakage in the pre-training data since all bugs are much more recent than Defects4J. Notably, GitBug-Java exclusively contains bugs from after the training data cutoff date of all models used in our experiments: CodeLlama-7b (September 2022) [24], gpt-4-0613 (September 2021)  [^2] https://platform.openai.com/docs/models, and gpt-3.5-turbo-0613 (September 2021) \@footnotemark.
@@
Fine-Tuning Hyperparameters We fine-tune CodeLLaMA with LoRA for each of our curated code representations with the same hyper-parameter settings: we set the learning rate to 5e-4 with cosine decay, max input length to 1024333Although longer input lengths enable the model to process more context and, potentially, fix more bugs, about four times more GPU memory is required when doubling the input length., training epoch to 2, and batch size to 16 per GPU, and we use Adam_W as the optimizer. For LoRA, we use a rank of 8, alpha of 16, dropout of 0.05, and inject the adaptation matrices in the q_proj and v_proj layers. Using the same hyper-parameter settings for each code representation ensures fair comparison. Each fine-tuning run is executed on a server with 4xA100 40GB GPUs.
@@
Inference Setup In inference, we employ beam search as our decoding strategy with a beam size of 10 per previous research [8]. Hence, for each bug, we generate 10 candidate patches. We use the HuggingFace transformers library to implement all fine-tuning and inference experiments. Inference is run on a single A100 40GB GPU.
3.3 Patch Assessment
Patch assessment is a notoriously hard task. In our evaluation, we compute the best metrics for that task [11, 8, 26]: 1) A plausible patch is defined as one that successfully passes all test cases. 2) An exact-match patch is textually identical to a developer-provided reference patch, incl. spaces and formatting. 3) An AST-match patch has an AST which is equivalent to the AST of the developer-provided reference patch. 4) A semantic-match patch is a patch deemed equivalent after manual assessment by an expert.
Plausible and exact-match patches are straightforward. Let us dwell on the two other kinds.
@@ 3.3 Patch Assessment
The major advantage of an AST-match patch is to compute performance regardless of formatting and indentation changes. It is also more scalable than manually checking patches for correctness without expertise in the programs under repair. The AST-match process involves converting plausible and reference patches into abstract syntax trees [36] and subsequently utilizing AST differencing [37] to compare their ASTs for discrepancies.
A semantic-match patch is the most costly assessment to get. In this paper, to assess semantic equivalence, the two first authors independently label all plausible but not AST/Exact match patches in a first round. For the patches the two first authors disagree upon, the third author breaks the tie.
For all four metrics, the higher the metric, the better the performance. We validate the candidate patches on a workstation with an 18-core Intel Core i9-10980XE CPU and 128 GB of RAM, operating under Ubuntu 22.04.3 LTS.
3.4 Methodology for RQ1
@@
The objective of RQ1 is to investigate the most effective code representations for fine-tuning an LLM for program repair. While existing research has delved into the utility of LLMs for program repair, the impact of the code representations, such as their realism, has been overlooked. It is known that variations in code representations may yield substantial differences in performance for fine-tuned LLMs [25]. Consequently, in RQ1, we empirically evaluate 6 realistic code representation pairs presented in 2.4 and measure their performance.
We fine-tune an LLM as described in 3.2. We prompt the model to generate 10 patches for each bug using beam search decoding. We then evaluate the generated patches as outlined in subsection 3.3, to measure the effectiveness of each code representation. We prompt the non fine-tuned CodeLLaMA-7B as a baseline.
We assess the statistical significance of performance difference by employing the McNemar test for each pairwise combination of representations. In our context, each statistical test looks at the binary outcomes of two representations (or models) evaluated on the same set of benchmark examples, according to semantical match. The null hypothesis ($H_{0}$) for the McNemar’s tests is that the distributions are indistinguishable, i.e. the row and column marginal frequencies in the 2x2 contingency table are equal, i.e., $H_{0}:P(A=1,B=0)=P(A=0,B=1)$, where A and B represent whether A and B, respectively, fixed a given bug.
3.5 Methodology for RQ2
@@
The objective of RQ2 is to evaluate the respective effectiveness of parameter-efficient and full-parameter fine-tuning. Generally, parameter-efficient fine-tuning methods represent a trade-off between computational cost and model performance, in order to train LLMs with limited computational resources. While traditional full-parameter fine-tuning approaches often yield better results, it comes at the expense of significantly higher memory requirements and a larger-scale fine-tuning dataset. In other words, fully fine-tuning an LLM on a small fine-tuning dataset may result in overfitting. In this experiment, we explore and compare the effectiveness of parameter-efficient and full-parameter fine-tuning in the specific context of program repair, which has never been done to the best of our knowledge.
@@
Baseline. We consider four baselines in RQ2. One is the baseline considered in RQ1. The second one comprises the full-parameter fine-tuned version of RepairLLaMA, a powerful but naive approach to fine-tuning a large language model for program repair. Here, we use the same hyper-parameters as in LoRA fine-tuning, apart from a lower learning rate of 2e-5. We employ the same statistical significance test as described in subsection 3.4 when comparing RepairLlama against these two baselines.
The third is Jiang et al.’s work [8]. We compare against the best performing model reported, the fine-tuned version of Incoder-6B [38]. The last one is RAP-Gen [10], a state-of-the-art fine-tuned LLM for program repair, which uses fix patterns retrieved from a codebase of previous bug-fix pairs to augment the buggy input.
3.6 Methodology for RQ3
Figure 4: The prompt used to prompt GPT-3.5 and GPT-4 as a strong baseline to generate patches.
@@ 3.6 Methodology for RQ3
Recently, related work [14, 39] has shown that GPT-3.5-Turbo and GPT-4 achieve state-of-the-art results on program repair. The objective of RQ3 is to study how RepairLLaMA compares against state-of-the-art ChatGPT-based program repair.
First, we zero-shot prompt gpt-3.5-turbo-0613 and gpt-4-0613 to generate 10 patches for each bug. The prompt is shown in Figure 4, it is built by integrating the best code representation with an effective zero-shot prompt from Zhang et al. [39]. The prompt instructs the LLM to generate the fixed code chunk to replace the <FILL_ME> token. Note that the same information signals are provided to both RepairLLaMA models and the GPT models, allowing for a fair comparison of the repair capabilities of the different models. This includes providing the same source code context, as well as the same fault localization information as input. We employ the same statistical significance test as described in subsection 3.4. We utilize OpenAI’s official APIs to call gpt-3.5-turbo-0613 and gpt-4-0613 to conduct our experiments on Dec. 1, 2023, except for the GitBug-Java experiments which are conducted on gpt-3.5-turbo-0125 on Nov. 17, 2024.
4 Experimental Results
4.1 Results of RQ1 (Code Representations for Fine-Tuning)
@@
Table II: Repair results of different code representations for fine-tuning an LLM for program repair (RQ1). Our best model, RepairLLaMA using IR4xOR2, significantly improves over the baseline in all test benchmarks.
{tblr}
row2 = c, column2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=4c, cell17 = c=4c, cell111 = c=4c, cell33 = r, cell34 = r, cell35 = r, cell36 = r, cell37 = r, cell38 = r, cell39 = r, cell310 = r, cell311 = r, cell312 = r, cell313 = r, cell314 = r, cell43 = r, cell44 = r, cell45 = r, cell46 = r, cell47 = r, cell48 = r, cell49 = r, cell410 = r, cell411 = r, cell412 = r, cell413 = r, cell414 = r, cell53 = r, cell54 = r, cell55 = r, cell56 = r, cell57 = r, cell58 = r, cell59 = r, cell510 = r, cell511 = r, cell512 = r, cell513 = r, cell514 = r, cell63 = r, cell64 = r, cell65 = r, cell66 = r, cell67 = r, cell68 = r, cell69 = r, cell610 = r, cell611 = r, cell612 = r, cell613 = r, cell614 = r, cell73 = r, cell74 = r, cell75 = r, cell76 = r, cell77 = r, cell78 = r, cell79 = r, cell710 = r, cell711 = r, cell712 = r, cell713 = r, cell714 = r, cell83 = r, cell84 = r, cell85 = r, cell86 = r, cell87 = r, cell88 = r, cell89 = r, cell810 = r, cell811 = r, cell812 = r, cell813 = r, cell814 = r, cell93 = r, cell94 = r, cell95 = r, cell96 = r, cell97 = r, cell98 = r, cell99 = r, cell910 = r, cell911 = r, cell912 = r, cell913 = r, cell914 = r, cell103 = r, cell104 = r, cell105 = r, cell106 = r, cell107 = r, cell108 = r, cell109 = r, cell1010 = r, cell1011 = r, cell1012 = r, cell1013 = r, cell1014 = r, vline2-15 = 1, vline4-15 = 2, vline2-15 = 2-10, hline1,3-11 = 2-14, hline2 = 3-14, & Code Representations Defects4J v2
(488 bugs)
@@
(162 bugs)
GitBug-Java
(90 bugs)
Plausible Exact
Match AST
Match Semantic
Match Plausible Exact
Match AST
Match Semantic
Match Plausible Exact
Match AST
Match Semantic
Match
1 IR3 x OR2 (no fine-tuning) 131 52 70 83 107 71 81 103 17 8 9 12
2 IR4 x OR2 (no fine-tuning) 107 50 60 69 95 65 72 91 19 11 12 13
3 IR1 x OR1 79 29 31 45 78 52 54 72 10 4 4 4
4 IR1 x OR3 41 15 17 24 39 21 21 37 6 1 1 1
5 IR1 x OR4 12 2 2 3 5 2 2 4 1 1 1 1
6 IR2 x OR2 198 121 122 139 118 69 77 108 23 16 17 19
7 IR3 x OR2 153 83 86 102 103 63 68 99 21 12 12 13
8 IR4 x OR2 (RepairLLaMA) 195 124 125 144 118 75 82 109 25 15 16 20
@@
In RQ1, we investigate the most effective code representations for fine-tuning an LLM for program repair. The results of the evaluation are presented in Table II, which shows the effectiveness of each code representation setting on both test benchmarks. The table is structured as follows: the first column displays the code representations, the second to fourth meta-columns show the repair effectiveness results on, respectively, the single-function subsets of Defects4J v2, HumanEval-Java, and GitBug-Java. Recall that the repair effectiveness evaluation is measured by the four patch assessment metrics described in subsection 3.3.
Our baselines (rows 1 and 2) use the original non-fine-tuned CodeLLaMA-7B, with the IR3xOR2 and IR4xOR2 code representations respectively. We computed the baseline performance for other representations, and these are the most effective way to prompt the non-fine-tuned model [8], making them the strongest possible baselines.
@@
Our results show that the IR3xOR2 baseline plausibly repairs 131 Defects4J v2, 107 HumanEval-Java bugs, and 17 GitBug-Java bugs. Moreover, it correctly repairs 52 Defects4J v2, 71 HumanEval-Java, and 8 GitBug-Java bugs with patches that textually exactly match the developer-written ones. Furthermore, when considering AST match, it can repair 70 Defects4J v2, 81 HumanEval-Java, and 9 GitBug-Java bugs with patches that are syntactically equal to the developer-written ones. Finally, when considering semantic match, the baseline can repair 83 Defects4J v2, 103 HumanEval-Java, and 12 GitBug-Java bugs with patches that are semantically equivalent to the developer-written ones. Those results show that CodeLLaMA-7B with the appropriate prompting is already a very strong model for program repair.
The difference between the first baseline (row 1, IR3xOR2) and the second baseline (row 2, IR4xOR2) is that IR4 includes the buggy code as a comment. Despite this extra signal, the baseline model cannot make use of it effectively. Indeed, it achieves worse performance, correctly repairing 14 (69 vs. 83 of IR3xOR2) less Defects4J bugs, 12 (91 vs. 103) less HumanEval-Java bugs, and 1 (13 vs. 12) more GitBug-Java bug. This finding clearly shows that without any fine-tuning, inserting extra information into the representation such as the buggy code can act as noise.
@@
The fine-tuned model’s effectiveness depends on the code representations. First, we observe that the three code representation pairs that do not have access to fault localization (IR1xORX, rows 3 to 5) perform considerably worse than both the baseline and other code representations. These results show that fault localization signals are crucial for program repair. To that extent, all representations that simply use the full function as the input and ask to “fix the bug”, can be considered too naive. Our work demonstrates that tailoring code representations with fault localization is a necessary step in the context of program repair. Recall that no pre-training objective has access to fault-localization signals, which means that fine-tuning is essential to build knowledge of fault-localization signals.
@@
Second, we observe that, within code representations that use fault localization signals (IR2, IR3, and IR4, rows 6 to 8), fine-tuned models significantly outperform the baseline on Defects4J v2 and GitBug-Java compared to the baselines (rows 1 and 2). Again, this clearly demonstrates the importance of fault localization signals in the input representation.
The only exception to this is in HumanEval-Java, where the fine-tuned IR3xOR2 model performs worse than the baseline IR3xOR2 model. We hypothesize that this phenomenon occurs due to catastrophic forgetting: fine-tuning improves the model for a downstream task at the expense of some loss of knowledge from previous training. In our case, the catastrophic forgetting of HumanEval-Java is explained by the difference in nature between the fine-tuning dataset and the benchmark. Megadiff, the fine-tuning dataset, contains diffs sampled from Java projects hosted on GitHub. The nature of such projects is closer to the nature of Defects4J and GitBug-Java samples, as these are also collected from open-source Java repositories. HumanEval-Java, on the other hand, is manually engineered on top of simple coding tasks not commonly found on large open-source projects, and thus has a different distribution of bug and fix kinds. These results highlight the need for careful curation of fine-tuning datasets and benchmarks. Additionally, it also supports the effectiveness of IR4xOR2, which is able, to a certain degree, to alleviate catastrophic forgetting.
The best model, RepairLLaMA, fine-tuned with the code representation IR4xOR2, plausibly repairs 195 (+64) Defects4J bugs than the baseline, exactly repairs 124 (+72), syntactically correctly repairs 125 bugs (+55), and semantically correctly repairs 144 bugs (+61) bugs, respectively. This highest performance is due to the presence of the buggy code in the input, which gives the model unique ingredients for understanding the problem and generating the patch.
@@
To strengthen the external validity of our analysis beyond Defects4J, we perform the same experiment on HumanEval-Java and GitBug-Java. On HumanEval-Java, RepairLLaMA also achieves better repair effectiveness than the baseline models (75 vs. 71 exact match, 82 vs. 81 AST match, and 109 vs. 103 semantic match). On GitBug-Java, RepairLLaMA also beats the baseline models (15 vs. 11 exact match, 16 vs. 12 AST match, and 20 vs. 13 semantic match). These results are consistent with the observations made on Defects4J.
Now, we discuss the statistical significance of our results. All exact p-values are provided as supplemental material [40]. We observe that there is a statistically significant difference between the golden model RepairLLaMA (IR4xOR2) and all other models, with very low p-values ($p\leq 1.27e-04$), with the exception of IR2xOR2 for which the test fails to reject the null hypothesis, this case is further discussed below. These results completely validate the importance of the task-specific signals embedded in our best-performing code representations.
Third, we discuss the impact of fine-tuning in aligning the repair-specific code representation and the pre-training objectives. IR2 and IR4 are input representations that include the same signals: the original function, fault localization, and the original buggy code. The key difference is that IR4 uses the same infilling scheme seen during pre-training, while providing the original buggy code in comments, while IR2 places two comments as delimiters of the buggy region. We see a marginal performance gap between both representations (e.g., 144 vs. 139 correctly fixed Defects4J bugs) that is not statistically significant ($p\approx 0.545$) [40]. We hypothesize that, while IR2 does not implicitly use the infilling scheme, fine-tuning effectively maps the delimiting comments of IR2 to the <FILL_ME> token used by IR4 and the pre-trained infilling scheme.
@@
Figure 5: Exact match patch generated by RepairLLaMA for Math-86 from Defects4J v2. In this multi-location bug, RepairLLaMA is able to fix two distant buggy locations.
@@
In addition to comparing performance, we explore the effectiveness of region-based representations in addressing multi-location bugs. Our findings reveal that RepairLLaMA correctly repairs 51 instances of such bugs. For example, RepairLLaMA correctly fixes a complex multi-chunk bug, Math-86, from the Defects4J v2, as illustrated in Figure 5. Math-86 presents two error sections that require simultaneous attention and correction: 1) the removal of an if block that throws an exception, and 2) the introduction of a new if condition. Note that these two sections have more than 20 lines of distance between each other, showing that RepairLLaMA can fix bugs where the multiple edit locations are far away from each other. To summarize, our experiments give clear evidence that RepairLLaMA’s representation enables the repair of a wide range of multi-location bugs.
@@
Figure 6: Exact match patch generated by RepairLLaMA for STRONGEST_EXTENSION from HumanEval-Java. In this multi-location bug, RepairLLaMA modifies two if blocks and encapsulates two other statements in a new if block.